# CD-NGP: A Fast Scalable Continual Representation for Dynamic Scenes

### v3

@@ 3.4 Ablation studies
Table 3: Comparisons between our CD-NGP and baselines on the long video sequences in our dataset. denotes the method trained offline. Metrics are measured on the first 1200 frames. The compared HexPlane baselines are trained with a chunk size of 300 frames.
@@
HexPlane-PR 20.90 0.286 1272MB 8 hours 12GB
HexPlane-GR 24.82 0.282 1272MB 8 hours 12GB
HexPlane 28.32 0.117 800MB 8 hours 74GB
StreamRF 24.99 0.266 13GB 4 hours 12GB
INV 27.07 0.086 1344MB 160 hours 12GB
Ours 27.15 0.190 201MB 10 hours 12GB
@@
[…]
CD-NGP 21M 30.23 0.198 75min 113MB
3D+4D 23M 29.34 0.243 82min 122MB
Voxel-C 21M 30.52 0.174 80min 114MB
Plane 14M 29.16 0.215 77min 83MB
MERF 7.5M 29.93 0.201 109min 57MB
MERF-C 7.5M 30.22 0.174 115min 57MB
3.4.2 Spatial representations and scalability analysis.
The proposed hash-based pipeline is not only practical with voxel hash grids but also compatible with other hash-based spatial backbones. To prove the flexibility and scalability of our method, we provide ablation studies on different spatial representations. We adopt the Eq. 5 to the MERF representation [38], defined as:
We provide quantitative results on plane representation like [17, 21] and hybrid representation (MERF) [38] on the DyNeRF dataset in Table 5. Since the element-wise product for plane representation in [17] produces unsatisfactory results due to the hash collision, we remove the $P_{m}(x,y,z)$ in Eq. 12 and fuse the features $Q_{mij}(i,j)$ by concatenation in the implementation. The 3D+4D representation similar to [34] is yielded by concatenating the spatial $(x,y,z)$ features and spatial-temporal $(x,y,z,t)$ features encoded by a 4D extension of Eq. 5. By comparing our CD-NGP with the naive 3D+4D representation, we show the advantage of our asymmetric encoding of spatial-temporal $(x,y,z,t)$ features explained in Eqs. 3 and 6: representing the features of spatial coordinates $(x,y,z)$ and temporal coordinate $(t)$ separately avoids feature collisions and lead to better results.
Table 5 also shows metrics of different feature fusion methods. We replace the element-wise sum in Eqs. 6 and 12 with vector concatenation and denote them as Voxel-C and MERF-C in Table 5. The hybrid representations (MERF and MERF-C) achieve comparable model sizes ($\sim$60MB) to methods for static scenarios [31, 21] without significant quality loss, reducing the average bandwidth cost to $0.2$MB/frame. Benefiting from larger feature dimensions, the vector concatenation leads to slightly better quality (especially when measured in LPIPS) by more training time, and the reasonable trade-off also shows the robustness of our method. Compared with the hash table designs in Table 4, the Voxel-C representation in Table 5 almost performs as well as the bulky $(18,18)$ configuration in Table 4 and reduces the model size by $80\%$. We list the per-scene results of the proposed continual learning method with different spatial representations on the DyNeRF dataset in Table 6.
Model Flame Salmon Cook Spinach
[…]
CD-NGP 26.37 0.04586 0.2886 30.76 0.02440 0.1793
Voxel-C 26.65 0.04243 0.2680 31.03 0.02288 0.1593
Plane 25.48 0.06058 0.3173 29.41 0.03487 0.2010
MERF 26.15 0.04952 0.2921 30.35 0.02931 0.1858
MERF-C 26.56 0.04435 0.2546 30.69 0.02609 0.1583
Model Flame Steak Sear Steak
[…]
CD-NGP 31.27 0.02132 0.1736 31.71 0.01917 0.1701
Voxel-C 31.64 0.01823 0.1403 32.02 0.01698 0.1424
Plane 30.26 0.02949 0.1804 30.56 0.02723 0.1811
MERF 31.01 0.02508 0.1719 31.44 0.02181 0.1698
MERF-C 31.13 0.02424 0.1506 31.66 0.01989 0.1413
Model Cut Roasted Beef Average
[…]
CD-NGP 31.03 0.02366 0.1801 30.228 0.026882 0.19834
Voxel-C 31.26 0.02205 0.1578 30.520 0.024514 0.17356
Plane 30.11 0.03302 0.1951 29.164 0.037038 0.21498
MERF 30.71 0.02765 0.1856 29.932 0.030674 0.20104
MERF-C 31.04 0.02450 0.1632 30.216 0.027814 0.17360
The detailed parameter numbers and bandwidth consumption are thus listed in Table 7. 3D Enc. and 2D Enc. denote the spatial encodings along voxels and planes respectively. The tuples $(M_{1},M_{2})$ in the cells denote the parameter numbers of the base encoder and each of the auxiliary encoders. Following [27, 43, 1, 48] we also report the minimum per-frame bandwidth consumption $\mathbf{B_{min}}$, which only depends on the parameter numbers of the auxiliary branches and is thus significantly lower than the average bandwidth consumption $\mathbf{B_{avg}}$. Since the base encoder and the density grid could be shared among different branches and transmitted only once in the beginning, the minimum bandwidth consumption is reduced to $0.065\sim 0.16$ MB per frame, which is 100×100$\times$100 × less than the voxel-based StreamRF [27] and 7×7$\times$7 × less than the pure MLP representation INV [48].
CD-NGP $(9.9M,0.37M)$ $0.042M$ $0.39M$ $21M$ 0.16 0.38
Voxel-C $(9.9M,0.37M)$ $0.046M$ $0.40M$ $21M$ 0.16 0.38
Plane - $(1.9M,0.082M)$ $0.048M$ $0.28M$ $14M$ 0.11 0.28
MERF $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.042M$ $0.16M$ $7.5M$ 0.061 0.19
MERF-C $(1.3M,0.049M)$ $(0.56M,0.023M)$ $0.052M$ $0.17M$ $7.5M$ 0.065 0.19
@@
We also conduct ablation studies on the field composition methods and show the results in Tables 9 and 10. The field composition process is defined as:
@@
Full composition in Table 9 denotes the naive composition of 2 fields from the base branch and the current branch respectively, i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ are the field from the base branch and $\sigma_{a}$ and $\mathbf{c}_{a}$ in Eq. 13 are the field from the current branch. Both branches are encoded with the same architecture as Eq. 3. Static-dynamic in Table 9 denotes a naive composition of the field from a static branch by the base encoder and a dynamic branch by the auxiliary encoders similar to [45], i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ in Eq. 13 are encoded without the $\Gamma(t)$ in Eq. 3. The online training strategies remain the same. Our CD-NGP outperforms the other methods in all metrics. Furthermore, as compared in Fig. 10, the compared composition methods in Figs. 10(c) and 10(d) suffer from salt and pepper noise because of the explicit composition, while our CD-NGP circumvents the problem by fusing features in the latent space and produces better results.

### v4

@@ 3.4 Ablation studies
Figure 10: PSNR curves of the first 2400 frames of seminar scene in our long video dataset.
Figure 11: Comparison of the model sizes of our method and other online baselines. Our method show superior scalability over the compared baselines.
@@
Table 6: Quantitative comparisons between our CD-NGP and baselines on the long video sequences in our dataset. Metrics are measured on the first 1200 frames and averaged among the different scenes. The compared HexPlane baselines are trained with a chunk size of 300 frames.
@@
HexPlane $\checkmark$ 28.32 0.117 800MB 8 hours 74GB
HexPlane-PR $\checkmark$ 20.90 0.286 1272MB 8 hours 12GB
HexPlane-GR $\checkmark$ 24.82 0.282 1272MB 8 hours 12GB
StreamRF $\checkmark$ 24.99 0.266 13GB 4 hours 12GB
INV $\checkmark$ 27.07 0.086 1344MB 160 hours 12GB
CD-NGP (Ours) $\checkmark$ 27.15 0.190 201MB 10 hours 12GB
@@
We also conduct ablation studies on the field composition methods and show the results in Tables 9 and 12. The field composition process is defined as:
@@
Full composition in Table 9 denotes the naive composition of 2 fields from the base branch and the current branch respectively, i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ are the field from the base branch and $\sigma_{a}$ and $\mathbf{c}_{a}$ in Eq. 13 are the field from the current branch. Both branches are encoded with the same architecture as Eq. 3. Static-dynamic in Table 9 denotes a naive composition of the field from a static branch by the base encoder and a dynamic branch by the auxiliary encoders similar to [45], i.e. $\sigma_{i}$ and $\mathbf{c}_{i}$ in Eq. 13 are encoded without the $\Gamma(t)$ in Eq. 3. The online training strategies remain the same. Our CD-NGP outperforms the other methods in all metrics. Furthermore, as compared in Fig. 12, the compared composition methods in Figs. 12(c) and 12(d) suffer from salt and pepper noise because of the explicit composition, while our CD-NGP circumvents the problem by fusing features in the latent space and produces better results.